OpenStudy (beccaboo333):

Subtract. (6x3 - 10x2 + 6x - 1) - (4x3 - 2x2 - x + 3)

OpenStudy (beccaboo333):

A. 2x3 - 12x2 + 5x + 2 B. 2x3 - 8x2 + 5x + 2 C. 2x3 - 8x2 + 7x - 4 D. 2x3 - 12x2 + 7x - 4

OpenStudy (solomonzelman):

\[(6x^3 - 10x^2 + 6x - 1) - (4x^3 - 2x^2 - x + 3)\]\[(6x^3 - 10x^2 + 6x - 1) + \color{red} {(-1) \times (4x^3 - 2x^2 - x + 3)}\] See the trick in this question?

OpenStudy (anonymous):

it is d

OpenStudy (anonymous):

WAITTTT

OpenStudy (solomonzelman):

We are NOT giving out plain answers in math, this is unacceptable!

OpenStudy (anonymous):

That was wrong anyways lol I looked at the wrong one :p

OpenStudy (beccaboo333):

i still dont fully understand @SolomonZelman

OpenStudy (solomonzelman):

Ok, I'll show you.

OpenStudy (solomonzelman):

\[x-b~~=~~x\color{red} {+(-b)}~~~~which~~is~~also~~same~~as,~~~~x+\color{red} {(-1) \times b}\]

OpenStudy (solomonzelman):

First, do you agree with this?

OpenStudy (beccaboo333):

uhm sure.. just know i completely do not understand any of this

OpenStudy (beccaboo333):

sorry :(

OpenStudy (solomonzelman):

It's fine, lets do a different example together.

OpenStudy (beccaboo333):

@Mertsj

OpenStudy (beccaboo333):

help?

OpenStudy (solomonzelman):

\[\color{red} {(3x^2-2x)-(x^2+4)}\]\[3x^2-(x^2)~~~~~~~+~~~~~-2x-(+4)\]\[2x^2~~~~~+~~~~~-2x-4x\]\[2x^2~~~~~+~~~~~-6x\]\[2x^2-6x\] you should be able to understand this process.

OpenStudy (mertsj):

Do you know what like terms are?

OpenStudy (beccaboo333):

ones that are similar to one another and what isnt is different ?